# CroCoDai: A Stablecoin for Cross-Chain Commerce

### v1

@@ 3.5 System Requirements
    Full Backing. At any time $t\geq 0$, the ratio of the value of the collateral tokens to the total number of stablecoins in the system exceeds a safety threshold $\theta$, i.e.,
    $C^{*}_{t}/D^{*}_{t}>\theta.$ (2)
We assume that $\theta>1$, i.e., system failure occurs when the ratio of the collateral value to the debt gets sufficiently close enough to $1$. The motivation behind this requirement is as follows: the fundamental guarantee that underpins crypto-backed stablecoins is that in the event of a system shutdown, there is enough collateral in the vaults to refund each user who holds $h$ stablecoins with a basket of tokens whose value equals $h$ units of the pegged asset. If price fluctuations or an attack would cause this requirement to be violated, then we can reasonably expect the stablecoin to lose its peg. This would be detrimental to the user experience as their tokens held in escrow would depreciate, causing financial losses.
Due to random price fluctuations, we cannot guarantee with absolute certainty that (2) will hold in all circumstances. However, we use our analysis to show that for reasonable choices of $\theta$ and the liquidation ratio $\gamma$, the probability that (2) is violated after a medium-length period (i.e., one day) is very small for historical data. We do this both for portfolios that approximate the current state of the DSS (as of late April 2023), and portfolios that are optimized for minimal risk given a selection of possible single-chain or cross-chain tokens. In Section 5, we also investigate under which conditions (2) continues to hold in the event of a black swan event such as a sudden token price collapse or a 51% attack on a blockchain.
Furthermore, we require the following properties under the threat and token price model.
  1. 1.
    Safety. Users do not lose tokens during stablecoin transfers.
  2. 2.
    Liveness. Stablecoin transfers eventually happen.
  3. 3.
@@
More formally, for the last property we consider a minority of corrupted price oracles whose goal is to convince the protocol to accept a price $p^{*}$ instead of a true base price $p$ – safety of collateral then holds if their probability of success vanishes exponentially as $|p-p^{*}|$ increases (see also Theorem 5.3 in Section 5.1).
@@
4 CroCoDai: A Cross-Chain Stablecoin
Figure 1: Overview of the single-chain DSS.
In this section, we present CroCoDai, which extends the functionality of the single-chain DSS (depicted in Figure 1) to a multi-chain setting (depicted in Figure 2). Our main design goal is to have stablecoin tokens that are backed by collateral on multiple blockchains, and which can be moved securely and efficiently between the different chains. This is non-trivial as transactions between different blockchains are not atomic by default, so our design incorporates a secure cross-chain protocol. In the following, we will refer to blockchains that store tokens as coin chains (as per [27]). The communication between the different blockchains is managed by a relay chain, which may be a coin chain or a separate chain. The main difference between the single-chain DSS and CroCoDai is the inclusion of relay nodes – depicted with double circles in Figure 2 – which is a non-intrusive change as the other DSS contracts remain unchanged.
In the following, we first discuss in detail which components as discussed in Section 2.2 remain local to each chain and which require interaction with the relay chain. We then discuss the essential components of the smart contracts on the coin chains and the relay chain. Finally, we discuss two ways to implement the relay chain: one that relies on a quorum of trusted nodes, and one that uses an existing blockchain. We will present three implementations of the former in Section 6.
4.1 Overview
Local Operations. In CroCoDai, each coin chain keeps track of the balance of stablecoins in accounts on that chain. The stablecoins can be exchanged locally between accounts. Furthermore, each coin chain has its own vaults that allow users to create CDPs using tokens on that blockchain. Since CDPs are local, liquidation auctions and savings rate mechanisms are also processed locally. Finally, price feeds are still predominantly local, although price feeds from the relay chain can add an additional source of price information.
Cross-Chain Operations. The relay chain has four main functions in CroCoDai. The first is to enable cross-chain transactions. In a cross-chain transaction, stablecoins are burned on one coin chain and an equivalent number is minted on another chain. The second function is to add/remove addresses that are authorized to call security-critical functions such as mint or burn on another contract – we call such addresses wards as per DSS terminology. Periodically, smart contracts may be added or deprecated – e.g., 1. to add support for an emerging blockchain platform, 2. because a critical bug has been found in an existing smart contract, 3. because a chain has undergone a hard fork that interferes with the functioning of the contract, or 4. to implement a more efficient version of a component (e.g., upgrading the collateral auction mechanism from Cat.sol to Dog.sol in the DSS). The third function is to update the system parameters discussed in Section 2.2. The final function is to provide an additional source of token price information to vaults through relay chain oracles – this may particularly benefit smaller blockchains that lack a mature DeFi ecosystem.
The mechanism for cross-chain transfers is as follows: a node that holds stablecoins on a coin chain must create a transaction that stores the tokens to be transferred in escrow – furthermore, the transaction takes as input two strings to represent the target blockchain and the address on the target blockchain. This is then added to a list where it awaits commitment. The nodes in the relay network monitor the coin chains, and once they detect a cross-chain transfer request, they will sign a commit message. When the coin chain receives the commit message, it will commit the transaction and burn the tokens held in escrow. A commit message is also sent to the target blockchain, upon which it will mint an equivalent number of stablecoins. If the transaction is aborted for any reason, then the stablecoins in escrow are returned to the sender.
Figure 2: Overviews of CroCoDai, a multi-chain stablecoin system based on the DSS. The relays (double circle) are components that are not shared with the single-chain DSS.
4.2 Coin Chain Components
Each coin chain has the following components. Coin: implements a contract that is similar to Dai.sol in the DSS, and processes transfers of stablecoins between users on the coin chain. There is only one coin contract per chain. Vault: implements contracts that are similar to Vat.sol and Join.sol in the DSS, which allow users to create CDPs and withdraw/deposit collateral and stablecoins. There can be multiple vaults per chain. Auction: implements auction contracts that liquidate the collateral of underwater CDPs, with any appropriate auction logic (not necessarily limited to Cat.sol or Dog.sol) in the DSS. There can be multiple auction contracts per chain. Savings: implements a savings rate contracts that are similar to Pot.sol and Jug.sol in the DSS. Bottom-Level Oracle: implements a price feed using a contract similar to Spot.sol in the DSS, combined with a Medianizer. Bottom-Level Relay: validates messages from the relay chain. It processes 1. adding wards to each of the aforementioned smart contracts, 2. incoming cross-chain transfers, 3. commitments of outgoing cross-chain transfers, and 4. updates of system parameters.
4.3 Relay Chain Components
The relay chain has the following components. Governance: allows users to transfer governance tokens. Users who hold governance tokens can vote on system-level actions such as adding wards, processing cross-chain transfers, and updating system-level parameters. Top-Level Oracle: The relay chain may implement system-level price feeds to complement the price feeds on the coin chains. Top-Level Relay: To process cross-chain transfers, the relay network must be able to read the state of the coin chains, in particular the list of cross-chain transfer requests.
4.4 Relay Chain Designs
In this section, we discuss two possible implementation designs of the relay chain. The first relies on a quorum of trusted nodes with known public keys – each trusted node runs an off-chain (full or light) blockchain client for each of the underlying chains to check the validity of cross-chain transfers, and the nodes maintain a distributed ledger for governance decisions. The second uses an existing blockchain, and uses smart contracts that implement the consensus logic of the underlying chains on-chain to verify cross-chain operations.
Off-Chain Blockchain Clients. In this approach, the stablecoin system is maintained by $n$ relay nodes with known public keys. The relay nodes use their (light) clients to detect and validate cross-chain transactions, and generate signatures using their private/public key pair to commit cross-chain transfers and approve governance decisions. In this setting, the only functionality of the bottom-chain relay contracts is to verify digital signatures. If different chains use different built-in signature schemes, or the same scheme with a different parametrization, then the relay nodes may need to generate multiple signatures for a single cross-chain transfer (this is safe if the threat model of Section 3.4 holds when each signature is generated). We discuss three different implementations of this approach in Section 6.
To ensure security, we note that it is vital that the relay nodes are controlled by independent entities: e.g., in the attack against the Ronin bridge, 5 out of 9 validators were successfully attacked [30], of which 4 belonged to the same entity [36].
On-Chain Consensus Logic. In this approach, the consensus logic of the coin chains is implemented in the top-level relay contracts, and the consensus logic of the relay chain is implemented in the bottom-level relay contracts. Essentially, the functionality of a light client – i.e., validating block headers and proving the state of a contract using the header – is implemented inside the smart contracts. This is the approach taken in, e.g., Xclaim [58]. For example, in Bitcoin’s Nakamoto Consensus [45] this means the ability to compute header hashes and compare them to the proof-of-work difficulty, whereas in Ethereum’s proof-of-stake it means validating that nodes who control a sufficient amount of stake have approved a block. Techniques to increase the efficiency of light clients have been proposed in the literature for both proof-of-work (e.g., Bitcoin [14]) and proof-of-stake blockchains (e.g., Algorand [42] and Mina [4]). Another variation of this approach is to encode the consensus logic of the blockchains in a zero-knowledge proof, e.g., using the zkBridge protocol [54]. Finally, the data on the different coin chains can be integrated using TrustBoost [52].
Although this design avoids a trust assumption on a quorum of relay nodes, current iterations of such designs have a large computational overhead, typically requiring several kilobytes of data sent to relay contracts and more than 10 seconds to generate signatures [42, 54]. As such, we focus on the off-chain design in Section 6.
5 Analysis
In this section, we analyze whether our system satisfies the requirements of Section 3.5. In Section 5.1, we focus on the technical properties of safety and liveness, full backing under the threat model, and safety of collateral. We then focus on full backing under the price fluctuation model: we present our approach in Section 5.2, and the evaluation using empirical data in Section 5.3. The proofs of the theorems in this section can be found in Appendix D.
Our analysis shows that CroCoDai is resilient to serious failures of multiple coin chains, as long as the over-collateralization of the CDPs, $\gamma$, is high compared to the total debt ceiling of the affected chains (which we will denote by $\zeta^{\prime}$). We emphasize that CroCoDai is therefore fundamentally more resilient than a single-chain stablecoin such as the DSS, which cannot tolerate the failure of its single coin chain. Our analysis also shows that a broader selection of collateral tokens leads to a lower probability of system failure than a single-chain system such as the DSS. Although the findings of our price analysis strongly depend on the choice of data, we emphasize that we included a period of notoriously high volatility (i.e., the FTX collapse) and that the risks would appear smaller if we used a less volatile period.
5.1 Safety Analysis
@@
Technical Safety and Liveness. For local operations, the consensus protocol of the coin chain on which the local transfer occurs ensures the safety and liveness of operations on this chain – see, e.g., [22]. This includes local stablecoin transfers, operations on collateral, auctions, and interest withdrawals. For cross-chain operations, safety is violated if tokens in a bottom-level relay are burnt without a corresponding amount being minted on the receiving chain, and liveness is violated if tokens are held in escrow forever. However, the relay chain satisfies safety and liveness in our threat model, i.e., the transfer is eventually committed and coins are burned and minted on the outgoing and incoming chains, respectively. An alternative way to reach this conclusion is to view cross-chain transfers in CroCoDai as simple cross-chain deals [27] with the sender as the only party who votes, and the relay chain as a certified blockchain – depending on the design (as discussed in Section 4.4), either $12f+$ 1 nodes or the smart contracts on the relay chain sign the commit vote.
@@
Sudden Crash of Tokens. Occasionally, the price of a cryptocurrency token may suddenly collapse, e.g., because of the discovery of a serious bug in one of its smart contracts, or because of off-chain events such as a bankruptcy or a legal inquiry into the organization responsible for its maintenance. As an example: the bankruptcy of Celsius caused the value of the CEL token to collapse, whereas the inquiry into FTX did the same for the FTT token and Serum (SRM). In this case, the collateral auctions are likely to fail to generate enough revenue and the system’s over-collateralization is under threat. In the following, we consider a worst-case scenario: the price of some tokens falling to zero within a single time slot, i.e., between a given time slot $t^{\prime}$ and time slot $t^{\prime\prime}$ – for simplicity, we assume $t^{\prime}=0$ and $t^{\prime\prime}=1$. We show that CroCoDai remains fully backed at time $t^{\prime\prime}$ if it is sufficiently over-collateralized and if the total debt created using tokens on the affected chains is limited.
Formally, let $M^{\prime}\subset[M]$ denote the set of affected tokens, and $\zeta^{\prime}=\sum_{m\in M^{\prime}}\zeta_{m}$ the total debt ceiling of these tokens. Let

### v4

@@ 3.4. System Requirements
    Sufficient Overcollateralization. At any time $t\geq 0$, the ratio of the value of the collateral tokens to the total number of stablecoins in the system exceeds a safety threshold $\theta>1$.
The motivation behind this requirement is as follows: if the value of the collateral becomes too low, then the system is no longer able to reimburse all stablecoin holders with an equivalently valued basket of tokens in case of system shutdown. As such, we can expect users to start selling their stablecoin tokens at a loss to redeem as much value as possible, causing the stablecoin to lose its peg. We assume that this sell-off may occur before the critical point is reached at which the value of issued stablecoins exceeds that of the collateral, because nervous investors may preempt the loss of the peg. As such, we choose a value of $\theta$ that is greater than $1$: a value of strictly $1$ corresponds to a stablecoin system that fails as soon as the aforementioned critical point is reached, whereas $\theta>1$ also covers scenarios in which system failure occurs earlier, in anticipation of the depegging.
Due to random price fluctuations, we cannot guarantee with absolute certainty that full backing will hold in all circumstances. However, we use our analysis to show that for reasonable choices of $\theta$ and the liquidation ratio, the probability that full backing is violated after a medium-length period (i.e., one day) is very small for historical data. In Section 4, we also investigate under which conditions full backing continues to hold in the event of a black swan event such as a sudden token price collapse or a 51% attack on a blockchain.
We also require the following properties under the threat and token price model.
  1. (1)
    Safety. Users do not lose tokens during stablecoin transfers and CDP operations.
  2. (2)
    Liveness. Stablecoin transfers and CDP operations eventually happen.
  3. (3)
@@
More formally, for the last property we consider a minority of corrupted price oracles whose goal is to convince the protocol to accept a price $p^{*}$ instead of a true base price $p$ – safety of collateral then holds if their probability of success vanishes exponentially as $|p-p^{*}|$ increases (see also Theorem 4.3 in Section 4).
@@
3.5. Illustrative Application
We consider an example of cross-chain commerce: a setting in which an asset – e.g., a basket of cryptocurrencies or an NFT – is auctioned on one chain, and for which there are interested bidders who own tokens on multiple other chains. For the duration of the auction, price stability is important because changes in the value of bids could lead to unnecessary resubmissions or withdrawals of bids, causing a diminished user experience and/or opportunity costs (xue2021hedging, ). The auction example allows us both to illustrate the end-to-end processes of creating and transferring stablecoins in CroCoDai, as well as to compare CroCoDai to existing approaches in terms of performance and user experience. For this example, we will refer to the chain on which the asset is sold as the ‘auction chain’ and to the chains on which bidders hold tokens as ‘coin chains’.
To achieve an auction with price stability, we consider two main approaches: 1. bidders use a centralized exchange (CEX) to swap their tokens into stablecoins on the auction chain, 2. bidders using CroCoDai’s built-in mechanism for cross-chain transfers. The first represents the most straightforward way to achieve a cross-chain auction in the current ecosystem by leveraging established exchanges such as Coinbase or Binance, whereas the second is our proposal. We also discuss other approaches at the end of this section.
Figure 4. Steps of a cross-chain exchange in the CEX baseline and CroCoDai. The edge labeling corresponds to the step numbers in Section 3.5.
Comparison. In the first approach, we assume the existence of a centralized exchange that can be used to buy and sell stablecoins on the coin chain and auction chain. In particular, the bidder would then complete the following steps:
  1. (1)
    Create an account with the CEX if she doesn’t have one already.
  2. (2)
    Transfer her tokens to the CEX to add it to her wallet on the CEX.
  3. (3)
    Sell the tokens in her account and buy stablecoins, or swap the tokens for stablecoins directly if supported by the CEX, and withdraw the stablecoins to her account on the auction chain.
  4. (4)
    Place a bid by transferring the stablecoins to the auction contract.
In CroCoDai, the bidder would complete the following steps:
  1. (1)
    Deposit her tokens into a CroCoDai vault.
  2. (2)
    Create new CroCoDai stablecoins and withdraw them to her account.
  3. (3)
    Issue request to the relay chain for a cross-chain transfer of her CroCoDai tokens on the coin chain to an account owned by her on the auction chain.
  4. (4)
    Place a bid by transferring the stablecoins to the auction contract.
The fundamental difference between the two approaches is the reliance on a trusted third party in the first approach. By contrast, the second approach relies on a decentralized relay chain. Furthermore, the CEX baseline is more likely to suffer from a lack of liquidity: whereas steps 1 and 2 in the first baseline always rely on the CEX as a counterparty to swap tokens, in CroCoDai the bidder can always create new stablecoins and transfer them to the auction chain via the relay chain. In terms of fees and transaction delays, there are no reasons to expect major differences between the two approaches, since the cost of sending and processing cross-chain messages is typically small compared to the cost of monitoring the supported coin chains through a blockchain node, and the latter is a requirement for all cross-chain systems.
Additional Intermediaries. In both approaches discussed above, either the CEX or CroCoDai’s relay chain acts an an intermediary to facilitate the cross-chain token transfer. Additional steps that involve intermediaries may be needed if 1) the CEX does not support sales of the bidder’s tokens, 2) CroCoDai does not accept the bidder’s tokens as collateral (e.g., because they belong to an unsupported type, or because the token type’s debt limit would be violated), 3) the auction contract does not accept any token sold by the CEX, or 4) the auction contract does not accept CroCoDai stablecoins. In those cases, a DEX may exists that allows for on-chain swaps between unsupported and supported tokens. The additional steps are involving DEX swaps are depicted in Figure 4 through step $2^{\prime}$, which replaces step 2, and step $4^{\prime}$, which replaces step $4$.
Other Approaches. A cross-chain auction can also be implemented using the framework of cross-chain deals (herlihy2021cross, ). However, this requires that the auctioneer creates smart contracts on each of the supported coin chains, instead of the just the auction chain. Finally, the bidder could create a wrapped version of her tokens on the coin chain, but this would typically require another trusted intermediary and is therefore similar to the CEX baseline.
3.6. Deployment
Real-world use of CroCoDai would require the deployment of the relay chain, the smart contracts (e.g., coin and vault contracts) on the supported smaller chains, and the development of tools that allow users to interact with the system. The costs of developing and deploying the relay chain are similar to those of existing bridges such as Celer [^6] https://cbridge.celer.network/bridge/ethereum-polygon/eth/ and Relay [^7] https://relay.link/bridge/ethereum. The deployment of smart contracts has roughly the same cost as in a fully siloed system, and are small over long run compared to fees (e.g., at a gas price of 20GWei and an ETH price of \$3000, deploying a 1m gas contract would cost \$60). The throughput and scalability of the relay chain are highly dependent on the design choices and implementation, e.g., Celer’s cBridge claims a latency of at most 20 minutes during peak demand periods (cbridge, ), whereas the zkBridge prototype has a latency of under 2 minutes (xie2022zkbridge, ) but at a high computational cost, while the Hephaestus protocol (belchior2023hephaestus, ) claims to be able to process 600 cross-chain transactions in under 6 seconds.
In practice, we expect the user interface design to be very similar to existing designs for the DSS: web apps such as summer.fi888 https://summer.fi/ allow users to interact with vaults and to process local and cross-chain transfers. The incentives for developers are the same as for Dai, i.e., a combination of fees, publicity (an app can draw attention to other services), and in some cases volunteer work. The active DAI ecosystem, including summer.fi, daistats.com (daistats, ), DeFiExplore, [^9] https://maker.defiexplore.com demonstrates that such incentives are sufficient in practice.
Finally, the system parameters in CroCoDai can be set to similar values as in the DSS. The overcollateralization rates, i.e., $\gamma$, will likely be in the order of $30-50\%$ for the supported token types, as they are in existing Dai vaults (daistats, ). Similarly, the debt limits for smaller blockchains are expected to be similar to those for ERC20 tokens such as Aave and Gnosis – as in the DSS, these are proposed and voted on by governance token holders.
4. Security Analysis
In this section, we analyze whether our system satisfies the following technical properties from Section 3.4: safety and liveness, full backing under the threat model, and safety of collateral. We first discuss our threat model and then present our results. In particular, we prove that if tokens whose total debt limit [^10] We note that price crashes in different tokens and platforms may occur simultaneously, e.g., the collapse of FTX also had consequences for other platforms such as Solana. is $\zeta^{\prime}$ suffer a price crash (Section 4.3) or platform-level attack (Section 4.4), then full backing continues to hold if $\theta<\gamma(1-\zeta^{\prime})$, where $\gamma$ denotes the liquidation ratio (see Section 2.2). In practice, this allows stablecoin holders to judge whether they find the risk of system failure low enough given their estimations of $\theta$ and $\zeta^{\prime}$. We conclude by comparing the security of CroCoDai to the two baselines from the introduction. The technical proofs of the main theorems can be found in Appendix A.
[…]
4.1. Security Model
In the following, we first discuss the threat assumptions that are specific to CroCoDai, and then discuss the threats and countermeasures shared with generic stablecoin and cross-chain systems.
In CroCoDai’s security model, a majority of the governance token holders are honest and online, and the relay chain satisfies safety and liveness. Each underlying blockchain is either compromised or not. If it is not compromised, then we assume that it guarantees safety and liveness for all of its transactions. However, both safety and liveness can be violated in a compromised blockchain – in particular, transactions in which collateral is sent to/from a CDP can be rolled back. We assume that the underlying communication networks satisfy partial synchrony, i.e., bounded delivery times, to ensure that nodes do not wait indefinitely for information to be transmitted between chains. It has shown that liveness cannot be guaranteed in the presence of even a single faulty node in a fully asynchronous model (fischer1985impossibility, ). We do note that CroCoDai itself provides synchrony to the overall cross-chain process by requiring that a transfer is completed on the sending chain before processing it on the receiving chain, as depicted in Figure 4b. For the price oracles, a majority is online and honest at all times, in which case they report a measurement error that is random and normally distributed. Furthermore, the output of each oracle is independent of the output of other oracles. Dishonest oracles can report arbitrary prices. We emphasize that these assumptions are common in the scientific literature, e.g., Byzantine fault tolerant protocols also require honest and online (super)majorities and partial synchrony (lamport1982byzantine, ).
As a cross-chain stablecoin system, CroCoDai also inherits threats that affect generic blockchains, cross-chain infrastructures, and stablecoins. Recent surveys that investigate the security of such systems are available for individual blockchains (homoliak2020security, ; zhang2019security, ), cross-chain systems (belchior2021survey, ; zamyatin2021sok, ; herlihy2021cross, ), and stablecoins (mell2023understanding, ). Threats that affect individual blockchains include the following categories:
  • Network-layer threats, which may prevent users from connecting with honest blockchain nodes, potentially violating safety (if the remaining nodes are malicious) or liveness. Examples include eclipse attacks on blockchain peers, DNS/routing attacks, and Denial-of-Service (DoS) attacks. Countermeasures include generic network security proposals such as DNSSEC and BGPsec, or blockchain-specific defences such as SABRE (apostolaki2019sabre, ).
    Consensus protocols that announce a single block proposer beforehand – e.g., many Proof-of-Stake protocols (brown2019formal, ) – may be vulnerable to targeted DoS attacks against the block proposer, violating liveness. CroCoDai itself does not introduce such a vulnerability because cross-chain transfers can be initiated by any of the relay nodes.
  • Consensus-layer threats, which may cause established transactions to be reverted or the network to accept incorrectly formed blocks. Examples include majority (51%) attacks or selfish mining. Countermeasures include protocol-level defences, e.g., finality/checkpointing protocols, and robust incentive mechanism design.
    Finality protocols are also helpful to CroCoDai implementations, because CroCoDai needs a logic for each supported chain to determine when a cross-chain transfer has been processed (i.e., highly unlikely to be reverted). Finality protocols provide such a logic by default.
  • Contract-layer threats, which may allow attackers or malicious owners to subvert the intended functionality of a smart contract. Examples include smart contract bugs or deliberate backdoors. Countermeasures include static code analysis or formal verification, potentially run by third parties in the case of malicious owners or insider threats. CroCoDai’s similarity to the DSS suggests that we can leverage a larger market for smart contract validation services.
  • Application-layer threats, e.g., bugs or backdoors in applications that support the stablecoin system, such as tools that allow users to interact with CDPs. As with contract-layer threats, countermeasures include static code analysis and formal verification.
The main threats to stablecoin systems as enumerated by researchers at NIST (mell2023understanding, ) are categorized as security, stability, and trust “issues”. The following security “issues” are mentioned: (1) Unauthorized or Arbitrary Minting of Stablecoins, (2) Collateral Theft, (3) Malicious Smart Contract Update and Hijack, (4) Data Oracles, (5) Exploiting the Underlying Blockchain, and (6) Writing Secure Software and Vulnerabilities. Of these threats, (1), (2), and (5) are covered by the analysis in Section 4.4, as all scenarios effectively result in threats to the system’s overcollaterization. Finally, (4) is covered by the analysis in Section 4.5, whereas (3) and (6) are generic blockchain threats as discussed previously, which can be addressed using the countermeasures mentioned in (homoliak2020security, ; zhang2019security, ).
Security threats to generic cross-chain systems are an active area of research (belchior2021survey, ; zamyatin2021sok, ), although some threats such as cross-chain replay attacks and data availability violations have been studied in the scientific literature. We discuss the fundamental properties of safety and liveness in CroCoDai in more detail below.
4.2. Technical Safety and Liveness
@@
For local operations, the consensus protocol of the coin chain on which the local transfer occurs ensures the safety and liveness of operations on this chain – see, e.g., (garay2015bitcoin, ) – under the assumption of partial synchrony. This includes local stablecoin transfers, CDP operations, auctions, and interest withdrawals. As such, tokens cannot be stolen during local transfers or CDP operations or locked forever in a CDP unless the coin chain is compromised. For cross-chain operations, safety is violated if tokens in a bottom-level relay are burnt without a corresponding amount being minted on the receiving chain, and liveness is violated if tokens are held in escrow forever. However, the relay chain satisfies safety and liveness in our threat model, i.e., the transfer is eventually committed and coins are burned and minted on the outgoing and incoming chains, respectively. An alternative way to reach this conclusion is to view cross-chain transfers in CroCoDai as simple cross-chain deals (herlihy2021cross, ) with the sender as the only party who votes, and the relay chain as a certified blockchain – depending on the design (as discussed in Section 3.3), either $12f+$ 1 nodes or the smart contracts on the relay chain sign the commit vote.
@@
4.3. Sudden Crash of Tokens
Occasionally, the price of a cryptocurrency token may suddenly collapse, e.g., because of the discovery of a serious bug in one of its smart contracts, or because of off-chain events such as a bankruptcy or a legal inquiry into the organization responsible for its maintenance. As an example: the bankruptcy of Celsius caused the value of the CEL token to collapse, whereas the inquiry into FTX did the same for the FTT token and Serum (SRM). In this case, the collateral auctions are likely to fail to generate enough revenue and the system’s over-collateralization is under threat. In the following, we consider a worst-case scenario: the price of some tokens falling to zero within a single time slot, i.e., between a given time slot $t^{\prime}$ and time slot $t^{\prime\prime}$ – for simplicity, we assume $t^{\prime}=0$ and $t^{\prime\prime}=1$. We show that CroCoDai remains fully backed at time $t^{\prime\prime}$ if it is sufficiently over-collateralized and if the total debt created using tokens on the affected chains is limited.
Formally, let $\mathcal{J}_{t}$ denote the set of all CDPs at time $t$, $c_{j\,t}$ the total number of collateral tokens held in CDP $j\in\mathcal{J}_{t}$ at time $t$, and $D_{j\,t}$ the outstanding debt resulting from Dai withdrawals/deposits from/to CDP $j\in J_{t}$ at time $t$. We denote the token type associated with CDP $j$ by $y_{j}\in M$. Let $C_{t}^{*}=\sum_{j\in J}p_{y_{j}\,t}c_{j\,t}$ denote the total value of the system’s collateral at time $t$ and $D_{t}^{*}=\sum_{j\in J}D_{j\,t}$ the total debt. Let $M^{\prime}\subset M$ denote the set of affected tokens and $\zeta^{\prime}=\sum_{m\in M^{\prime}}\zeta_{m}$ the total debt ceiling of the affected tokens. Let